# Assessing Historical Structural Oppression Worldwide via Rule-Guided Prompting of Large Language Models

### v1

@@ IV-B Main Results
TABLE I: Model performance under different prompting strategies.
@@
Figure 1: Pearson correlations between LLM predictions and human annotations across countries.
@@ IV-C Error Analysis
Figure 2: Number of responses by difference between LLM assignment and human assignment, grouped by model.
@@ IV-C1 Incorrect Estimation of Severity
Figure 3: Number of responses by difference between LLM assignment and human assignment, grouped by prompt.
@@ IV-C2 Reasons for Incorrect Estimations
Figure 4: Highest Divergence between Human and LLM ratings across all models and prompt strategies.
Among all of the top divergences in our experiments, the majority are either produced by GPT-3.5 Turbo, in the CoT setting, or are identities in the country of Alegria (Figure 4). One example of CoT prompting leading to misunderstanding in the provided response is shown in this assignment provided by Gemini 1.5 Pro: For response “Indian” in country “United States”, the human annotator assigned an oppression level of 2, with the assumption that the response refer to the identity of Americans of Indian origin. The model assigned this response a level of 5, with the assumption that the response refers to Native American, arguing that “… [Native Americans/Indigenous peoples in the United States] have experienced extreme forms of violence, exclusion, and marginalization, resulting in long-term structural disadvantage.” While the reasoning itself is valid, the divergence arises from the model’s incorrect assumption about which identity the response denotes.
@@ V Discussion
This study has limitations. Our current study only examines ethnicity-related identity, not other important aspects of identity (e.g., gender identity, sexual orientation, disability identity). Additionally, the annotated set (n=334) spans 10 countries and, while intentionally diverse, cannot represent the full global distribution of identities and contexts. The transformation of multidimensional oppression into a five-level ordinal scale necessarily abstracts complex realities; different rubrics or task decompositions (e.g., legal status, economic exclusion, exposure to violence) might yield different performance profiles. Although we train annotators, adjudicate disagreements, and document sources and rationales, expert judgment is itself contextual and may reflect scholarly debates or data gaps in particular regions or languages. Finally, we conduct all prompting in a standardized template; alternative designs (e.g., richer exemplars, multi-turn retrieval-augmented reasoning) could shift outcomes.
Overall, our findings highlight the importance of structured, theory-informed prompting in aligning LLM outputs with expert-annotated oppression scores. Reasoning under rubric-guided rules, rather than unguided Chain-of-Thought (CoT) reasoning, is most effective at reducing error and improving reliability. Despite these additions, systemic miscalibration exists, particularly for ambiguous identities and politically complex regions, reinforcing the need for calibration against human-annotated benchmarks.
@@ Acknowledgments
We used ChatGPT for language editing to improve clarity and assist with LaTeX formatting. The use of generative AI in this manuscript adheres to ethical guidelines for use and acknowledgment of generative AI in academic research, as outlined in Mann et al. [38]. This manuscript has been thoroughly vetted for accuracy, and we assume all responsibility for the integrity of this contribution.

### v2

@@ IV-B Main Results
TABLE I: Model performance under different prompting strategies. The best results for each model are highlighted in bold.
@@
Figure 2: Pearson correlations between LLM predictions and human annotations across countries. Brazil has the highest alignment at $r$ = 0.86 while Algeria and Madagascar have the lowest alignment around $r$ = 0.5.
@@ IV-C Error Analysis
Figure 3: Distribution of absolute differences between LLM-predicted oppression scores and human expert annotations across 334 identity-country pairs, grouped by model. GPT-3.5-Turbo produces the highest number of high-magnitude errors ($|\text{difference}|\geq 2$), while GPT-4o-mini produces the fewest deviations.
@@ IV-C1 Incorrect Estimation of Severity
Figure 4: Distribution of absolute differences between LLM-predicted oppression scores and human expert annotations across 334 identity-country pairs, grouped by prompt method. Chain-of-Thought (CoT) prompting produces the highest number of high-magnitude errors ($|\text{difference}|\geq 2$), while Rule-Guided prompting produces the fewest deviations.
@@ IV-C2 Reasons for Incorrect Estimations
Figure 5: Highest Divergence between Human and LLM ratings across all models and prompt strategies. Majority of divergences occur from identities in Algeria, or are produced from scores generated by either CoT prompting or GPT-3.5-turbo.
Among all of the top divergences in our experiments, the majority are either produced by GPT-3.5 Turbo, in the CoT setting, or are identities in the country of Algeria (Figure 5). One example of CoT prompting leading to misunderstanding in the provided response is shown in this assignment provided by Gemini 1.5 Pro: For response “Indian” in country “United States”, the human annotator assigned an oppression level of 2, with the assumption that the response refer to the identity of Americans of Indian origin. The model assigned this response a level of 5, with the assumption that the response refers to Native American, arguing that “… [Native Americans/Indigenous peoples in the United States] have experienced extreme forms of violence, exclusion, and marginalization, resulting in long-term structural disadvantage.” While the reasoning itself is valid, the divergence arises from the model’s incorrect assumption about which identity the response denotes.
@@ V Discussion
This study has limitations. Our current study only examines ethnicity-related identity, not other important aspects of identity (e.g., gender identity, sexual orientation, disability identity). Additionally, the annotated set (n=334) spans 10 countries and, while intentionally diverse, cannot represent the full global distribution of identities and contexts. Similarly, the sample sizes for certain countries are relatively small. While we report point estimates of model performance (e.g., Pearson’s $r$, Spearman’s $\rho$, Cohen’s $\kappa$), we do not include 95% confidence intervals around these statistics, which would more fully capture the uncertainty introduced by small group sizes. As such, these metrics should be interpreted as suggestive rather than definitive; future work with larger, more balanced datasets is needed to validate the observed patterns and assess the consistency of model performance across diverse subpopulations. The transformation of multidimensional oppression into a five-level ordinal scale necessarily abstracts complex realities; different rubrics or task decompositions (e.g., legal status, economic exclusion, exposure to violence) might yield different performance profiles. Although we train annotators, adjudicate disagreements, and document sources and rationales, expert judgment is itself contextual and may reflect scholarly debates or data gaps in particular regions or languages. We also did not compute intercoder agreement metrics such as Cohen’s $\kappa$ for the ground-truth annotations and instead used a consensus-based approach. This method should be complemented with formal reliability assessments in future work to strengthen reproducibility. Finally, we conduct all prompting in a standardized template; alternative designs (e.g., richer exemplars, multi-turn retrieval-augmented reasoning) could shift outcomes.
Overall, our findings highlight the importance of structured, theory-informed prompting in aligning LLM outputs with expert-annotated oppression scores. Reasoning under rubric-guided rules, rather than unguided Chain-of-Thought (CoT) reasoning, is most effective at reducing error and improving reliability. Despite these additions, systemic miscalibration exists, particularly for ambiguous identities and politically complex regions, reinforcing the need for calibration against human-annotated benchmarks. Future work could incorporate Retrieval-Augmented Generation (RAG) or few-shot calibration to reduce bias in underperforming regions. Integrating country-specific histories via dynamic context windows could further enhance cultural specificity of LLM judgements.
@@ Acknowledgments
We used ChatGPT for language editing to improve clarity and assist with LaTeX formatting. The use of generative AI in this manuscript adheres to ethical guidelines for use and acknowledgment of generative AI in academic research, as outlined in Mann et al. [porsdam2024guidelines]. This manuscript has been thoroughly vetted for accuracy, and we assume all responsibility for the integrity of this contribution.
-A Additional Details of Methods
-A1 Survey Question
We provide the question in several other languages below:
  • Italian: Come descriveresti le tue origini etniche?
  • French: Comment décririez-vous vos origines ethniques?
  • Spanish: ¿Cómo describiría su trasfondo étnico?
  • Mandarin: [Uncaptioned image]
  • Arabic: [Uncaptioned image]
  • Hindi: [Uncaptioned image]
  • American Sign Language (ASL): https://youtu.be/p8AJYSlubkM